# Deep Instruction Tuning for Segment Anything Model

### v1

@@ Abstract
Segment Anything Model (SAM) exhibits powerful yet versatile capabilities on (un) conditional image segmentation tasks recently. Although SAM can support various segmentation prompts, we note that, compared to point- and box-guided segmentation, it performs much worse on text-instructed tasks. We argue that deep text instruction tuning is key to mitigate such shortcoming caused by the shallow fusion scheme in its default light-weight mask decoder. In this paper, two deep instruction tuning (DIT) methods are proposed, one is end-to-end and the other is layer-wise. With these tuning methods, we can regard the image encoder of SAM as a stand-alone vision-language learner in contrast to building another deep fusion branch. Extensive experiments on three highly competitive benchmark datasets of referring image segmentation show that a simple end-to-end DIT improves SAM by a large margin, with layer-wise DIT further boosts the performance to state-of-the-art. Our code is anonymously released at: https://github.com/wysnzzzz/DIT.
Machine Learning, ICML
1 Introduction
Figure 1: The comparison among the default fusion scheme of SAM (with ViT-B) and our deep instruction tuning methods on RefCOCO (Yu et al., 2016), i.e., the end-to-end and layer-wise ones, respectively [^2] Three methods are trained using the frozen ViT-B encoder.. In SAM, the text prompts only interactive with the visual tokens with the shallow mask decoder, which is insufficient for cross-modal fusion and text-guided segmentation. Our simple deep tuning methods can improve the performance by a large margin without significantly changing the architecture of SAM.
@@ 1 Introduction
Figure 2: Illustrations of the default fusion of SAM (a) and our deep instruction tuning (DIT) methods (b), i.e., the end-to-end (red) and layer-wise (blue) ones. Compared with the default shallow fusion, DITs regard the visual encoder as an deep multi-modal learner for cross-modal fusion and visual feature learning, thereby achieving the full interactions between text and image for text-guided SAM.
@@
Motivated by the great success of large language models (Ouyang et al., 2022; Touvron et al., 2023), Kirillov et al. (2023) recently propose a Segment Anything project to build a foundation model for image segmentation, termed Segment Anything Model (SAM). SAM demonstrates an impressive yet versatile capacity on various image segmentation tasks (Long et al., 2015; He et al., 2017; Kirillov et al., 2019) by pre-training on over one billion masks. Moreover, it can support interactive segmentation conditioned on various input prompts, such as points, boxes or texts.
However, for SAM to follow text instructions, as shown in Fig. 2, it only achieves 55.7 accuracy when fine-tuned on RefCOCO (Yu et al., 2016), which is much inferior than following points or boxs (Kirillov et al., 2023). Text instructions naturally exhibit linguistic ambiguities, while location prompts (points and boxes) are precise. Even worse, the shallow fusion scheme in SAM further aggravates such weakness and degrades its multi-modal ability.
@@
More concretely, for either point, box or text, input prompt is first encoded by the prompt encoder of SAM, then it is directly fed into the light-weight encoder of SAM for mask prediction. In SAM, prompts can only interact to a limited extent with visual tokens via two cross-attention layers in its mask decoder (Kirillov et al., 2023). Although such mechanism is well suited for point and box prompts which are relatively easy to follow, it is insufficient for SAM to infer the intentions behind texts step-by-step. Oftentimes, SAM needs to resolve linguistic ambiguity first through the attributes and relationships of the referent, then the object can be masked using its powerful and class-agnostic segmentation ability. To this end, we argue that
@@
To validate this argument, we propose two deep instruction tuning (DIT) methods for SAM. Different from existing RIS models  (Huang et al., 2020; Deng et al., 2021; Kim et al., 2022; Li et al., 2022) that build deep fusion branches upon image encoders, we focus on improving text instruction following ability of SAM without modifying its structure. In particular, we draw inspiration from prompt tuning (Zhou et al., 2022a, b) and directly extend SAM to a deep vision-language model. Enabled by the flexibility of Transformer-based encoder (Dosovitskiy et al., 2020), we convert prompt words to input tokens and concatenate them with visual tokens for deep multi-modal fusion and feature learning, which forms the first DIT scheme in this paper, we termed end-to-end one (E-DIT). As shown in Fig. 2 and Fig. 2, this simple method improves SAM by a large margin, i.e., +11.8% on RefCOCO val. To further enhance instruction learning, we also propose a layer-wise instruction tuning method (L-DIT) that independently projects and inserts the text words into each layer of SAM. This layer-wise method has a similar principle with soft prompt tuning (Jia et al., 2022). The difference is that the inserted tokens are the projected text embeddings that vary across examples, instead of task-specific and parameterized prompt tokens.
With the proposed DIT methods, we apply SAM to the popular text-guided Referring Image Segmentation (RIS) (Liu et al., 2023d, b), and term the new models as DIT-SAMs. Extensive experiments are conducted on three highly competitive benchmarks, namely RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016) and RefCOCOg (Mao et al., 2016; Nagaraja et al., 2016). The experimental results not only show the great advantages of DIT-SAMs over SAM, e.g., up to 11.6% improvement on RefCOCO testA, but also confirm the competitiveness of DIT-SAM as a stand alone RIS model. Overall, we provide a fast and feasible solution to extend SAM’s modal-modal ability.
@@
  • We propose two simple yet effective deep instruction tuning (DIT) methods for SAM, which are end-to-end and layer-wise, respectively. These DIT methods provide feasible solutions to efficiently extend the multi-modal capabilities of SAM.
@@
  • The proposed DIT-SAMs not only greatly outperforms SAM on RIS benchmarks, but also shows great competitiveness against existing SOTA methods of RIS.
2 Related Work
2.1 Image Segmentation
@@
As a fundamental task in computer vision, image segmentation (Ronneberger et al., 2015; Long et al., 2015; He et al., 2017; Badrinarayanan et al., 2017) requires pixel-level understanding of a given image. Various segmentation-related tasks have been investigated, such as semantic segmentation (Long et al., 2015; Badrinarayanan et al., 2017; Chen et al., 2017; He et al., 2017; Xie et al., 2021), which categorizes individual pixels into a predetermined set of classes, and instance segmentation, primarily concerns itself with the recognition and delineation of distinct object instances (He et al., 2017; Tian et al., 2020; Wang et al., 2020), Panoptic segmentation, which merges semantic and instance segmentation tasks by assigning class labels and individual instance identification (Kirillov et al., 2019; Li et al., 2019). Recently, the Segment Anything project (Kirillov et al., 2023) proposes a powerful model for image segmentation, termed SAM. This project includes the creation of an unprecedentedly large segmentation dataset so SAM is trained with about 1 billion image-mask pairs. Due to this large-scale pre-training, SAM exhibits robust and versatile performance across various segmentation challenges. Notably, SAM is also capable of handling different types of prompts for conditional segmentation, such as point, box and text. However, as mentioned above, SAM works much worse with texts than the other types of prompts.
2.2 Referring Image Segmentation
@@
Referring image segmentation (RIS) is a text-conditioned segmentation task that grounds the referent with binary masks according to the given natural language expression (Hu et al., 2016). Early research in RIS (Yu et al., 2016; Liu et al., 2019) typically adopts a two-stage pipeline and frames RIS as a problem of region-text matching. One-stage RIS models (Liu et al., 2017; Li et al., 2018; Margffoy-Tuay et al., 2018) recently garner growing attention from both academia and industry. In these one-stage models, text features are often embedded into the vision network and directly output the mask of the referent. Inspired by the success of the Transformer, various Transformer-based RIS models (Zhu et al., 2022; Hu et al., 2023; Liu et al., 2023c, a; Wu et al., 2023a; Xu et al., 2023a) are recently proposed to improve the multi-modal reasoning ability for RIS. VLT (Ding et al., 2022) and SeqTR (Zhu et al., 2022) use a encoder-decoder framework to fuse visual and language inputs into the decoder. ReSRT (Kim et al., 2022) employs two separate Transformer encoders to handle two different modalities. LAVT (Yang et al., 2022) add a ladder-side language pathway for the cross-modal fusion of text features and visual tokens. UniLSeg (Liu et al., 2023f) employs symmetric and multi-modal Transformer networks for multi-modal fusion, while UNINEXT (Yan et al., 2023) utilizes a multi-modal Transformer encoder for the same purpose. LISA (Lai et al., 2023) and SEEM (Zou et al., 2024) focus on text-instructed open-world segmentation, they introduce much more training data and the use of giant LLMs (e.g., LLaMA-7B(Touvron et al., 2023)). Apart from network design, recent research  (Luo et al., 2020b; Li & Sigal, 2021; Liu et al., 2023f) also emphasizes early fusion to enhance performance.
2.3 Prompt Tuning
With the ever increasing parameter size of pre-trained models (Devlin et al., 2018; Dosovitskiy et al., 2020; Touvron et al., 2023), the direct fine-tune on downstream tasks becomes prohibitively expensive. In this case, prompt tuning (Gu et al., 2021; Liu et al., 2021; Wei et al., 2021; Han et al., 2022) is proposed to use hand-craft or learnable prompt tokens to mitigate distribution shifts between pre-training and downstream tasks. The significant achievements of prompt tuning in natural language processing (NLP) have also inspired its application in the fields of computer vision and vision-language research (Liu et al., 2023e). CoOp (Zhou et al., 2022b) remains CLIP parameters untouched and uses learnable parameters as soft prompts for the input text, thereby maximizing CLIP’s zero-shot retrieval capabilities. To address overfitting issue on base classes, CoCoOp (Zhou et al., 2022a) proposes instance adaptation by incorporating the visual features of each image into the learnable prompts. RPO (Lee et al., 2023) enhances the robustness of VL models by using masked attention to prevent the added learnable embeddings from altering the original internal representations of the model. More recently, Wu et al. (Wu et al., 2023b) propose a method of approximated prompt tuning to reduce the additional cost of adding prompt tokens.
@@ 2.3 Prompt Tuning
Figure 3: Overall pipeline of Deep Instruction Tuning (DIT). We propose (a) End-to-end DIT (E-DIT) that appends language instructions before visual features to achieve early-fusion for enhancing multi-modal reasoning capability; (b) Layer-wise DIT (L-DIT) that layer-wisely projects to interact with the existing sequences in the well learned semantic space.
3 Method
3.1 Problem definition
@@
Referring image segmentation aims to segment the referred object in an image $I$ conditioned on the paired language expression $T$, its objective is formulated as
$\arg\min_{\theta}\mathcal{L}(\mathcal{F}(I,T;\theta),M),$ (1)
@@
To achieve this task, SAM (Kirillov et al., 2023) first extracts vision and text features through its vision and language encoder, based on which the light-weight decoder takes as input to predict the binary masks. Thus, the model $\mathcal{F}$ instantiated by SAM is
$\mathcal{F}(I,T;\theta)=\mathcal{F}_{d}(\mathcal{F}_{v}(I),\mathcal{F}_{t}(T)),$ (2)
@@
where ${F}_{v}$, ${F}_{t}$ and ${F}_{d}$ refer to visual backbone, language encoder and mask decoder, respectively.
@@
The mask decoder is very lightweight and only contains two cross-attention layers, which are in charge of both mask decoding and multi-modal interactions. In this case, SAM can be regarded as one of the late-fusion models which proved to be less effective than the early-fusion ones (Luo et al., 2020a, b; Zhu et al., 2022; Liu et al., 2023d) in addressing the complicated multi-modal reasoning tasks. Meanwhile, due to a lack of training on cross-modal data, text-guided SAM performs worse than the point- or box-guided SAM, leading to much inferior cross-modal abilities than existing SOTA methods in RIS (Zhu et al., 2022; Liu et al., 2023d).
@@
We introduce Deep Instruction Tuning (DIT) for SAM to better follow the intentions reflected in the language expression $T$. In particular, DIT keeps the whole SAM architecture unchanged and directly uses the image encoder as a strong multi-modal fusion network. The text prompts can continually interact with the visual contents through the whole feature learning process, helping SAM progressively locate the referent. Specifically, SAM with DIT can be then formulated as
$\mathcal{F}(T,I;\theta)=\mathcal{F}_{d}(\mathcal{F}_{v}(I,\phi(\mathcal{F}_{t}(T))),\mathcal{F}_{t}(T)).$ (3)
@@
Here, $\phi(\cdot)$ is a mapping function to project the text features onto visual space.
3.2 End-to-end Instruction Tuning

### v2

@@ Abstract.
Recently, Segment Anything Model (SAM) has become a research hotspot in the fields of multimedia and computer vision, which exhibits powerful yet versatile capabilities on various (un) conditional image segmentation tasks. Although SAM can support different types of segmentation prompts, we note that, compared to point- and box-guided segmentations, it performs much worse on text-instructed tasks, e.g., referring image segmentation (RIS). In this paper, we argue that deep text instruction tuning is key to mitigate such shortcoming caused by the shallow fusion scheme in its default light-weight mask decoder. To address this issue, we propose two simple yet effective deep instruction tuning (DIT) methods for SAM, one is end-to-end and the other is layer-wise. With minimal modifications, DITs can directly transform the image encoder of SAM as a stand-alone vision-language learner in contrast to building another deep fusion branch, maximizing the benefit of its superior segmentation capability. Extensive experiments on three highly competitive benchmark datasets of RIS show that a simple end-to-end DIT can improve SAM by a large margin, while the layer-wise DIT can further boost the performance to state-of-the-art with much less data and training expenditures. Our code is released at: https://github.com/wysnzzzz/DIT.
footnotetext: †Corresponding Author.
1. Introduction
@@
Figure 1. Illustration of the default fusion of SAM (a) and our deep instruction tuning (DIT) methods (b), i.e., the end-to-end (red) and layer-wise (blue) ones. Compared with the default shallow fusion, DITs regard the visual encoder as an deep multi-modal learner for cross-modal fusion and visual feature learning, thereby achieving the full interactions between text and image for text-guided SAM.
@@
Motivated by the great success of large language models (LLMs) (Ouyang et al., 2022; Touvron et al., 2023), Kirillov et al. (Kirillov et al., 2023) recently propose a Segment Anything project to build a foundation model for image segmentation, termed Segment Anything Model (SAM). SAM demonstrates an impressive yet versatile segmentation capacity by pre-training on over one billion masks. Moreover, it can support interactive segmentation conditioned on various input prompts (Kirillov et al., 2023), such as points, boxes or texts. Its outstanding ability also arouses great interest from the communities of multimedia (Cheng et al., 2023a, b; Ye et al., 2023; Gao et al., 2023) and computer vision (Wu et al., 2023a; Deng et al., 2024; Cheng et al., 2024), and has been recently transferred to various segmentation tasks, such as medical image segmentation (Guo et al., 2024; Xie et al., 2024; Quan et al., 2024), urban image segmentation task (Zhang et al., 2024a), and marine animal segmentation (Zhang et al., 2024b), or to provide instance masks for image editing tasks (Xie et al., 2023; Gao et al., 2023).
However, the text-awareness of SAM is much inferior than following the prompts of points or boxes (Kirillov et al., 2023). For instance, on true-color segmentation task (Xiao et al., 2024), SAM can achieve state-of-the-art performance of 85.2% mIoU with box guidance. In stark contrast, SAM can only achieve 58.0% mIoU when fine-tuned on the RefCOCO benchmark (Yu et al., 2016) of the popular text instructed task, i.e., referring image segmentation (RIS) (Yu et al., 2016) as shown in Fig. 2. To explain, text instructions naturally exhibit linguistic ambiguities, while location prompts (points and boxes) are precise. Meanwhile, recent Transformer-based vision models (Dosovitskiy et al., 2020; Liu et al., 2021b) have well incorporated with this type of spatial information, e.g., position embedding, and it is natural for SAM to well master these location prompts. In addition, the collection of instance-wise text descriptions is more laborious, and even the public Segment Anything project also does not release the text-conditioned annotations (Kirillov et al., 2023).
@@
Even worse, the shallow fusion scheme in SAM further aggravates such weakness and degrades its multi-modal ability. Concretely, for either point, box or text, input prompt is first encoded by the prompt encoder of SAM, then it is directly fed into the light-weight encoder for mask prediction. In SAM, prompts can only interact to a limited extent with visual tokens via two cross-attention layers in its mask decoder (Kirillov et al., 2023). Although such mechanism is well suited for point and box prompts which are relatively easy to follow, it is insufficient for SAM to infer the intentions behind texts step-by-step. Oftentimes, SAM needs to resolve linguistic ambiguity first through the attributes and relationships of the referent, then the object can be segmented using its powerful and class-agnostic segmentation ability. To this end, we argue that
@@
To validate this argument, we propose two deep instruction tuning (DIT) methods for SAM. Different from existing RIS models  (Huang et al., 2020; Deng et al., 2021; Kim et al., 2022; Li et al., 2022) that build deep fusion branches upon image encoders, we focus on improving text instruction following ability of SAM without modifying its structure. This property can help DITs make full use of its superior segmentation capability while reducing the training cost to a large extent. In particular, we directly extend SAM to a deep vision-language model. Enabled by the flexibility of Transformer-based encoder (Dosovitskiy et al., 2020), we convert text instruction words to input tokens and concatenate them with the visual ones for deep multi-modal fusion and feature learning, which forms the first DIT scheme in this paper, we termed end-to-end DIT (E-DIT) , as depicted in Fig. 3. This simple method improves SAM by a large margin, e.g., +11.8% on RefCOCO val, as shown in Fig. 2. To further enhance instruction learning, we also propose a layer-wise instruction tuning method (L-DIT) that independently projects and inserts the text words into each layer of SAM. This layer-wise method has a similar principle with soft prompt tuning (Jia et al., 2022). The difference is that the inserted tokens are the projected text embeddings that vary across examples, instead of task-specific and parameterized prompt tokens that are static. Besides, DITs aim to project the text words onto the well-learned visual semantic space of SAM, while prompt tuning is often to help the model better adapt the downstream task without full tuning (Jia et al., 2022).
Figure 2. The comparison among the default fusion scheme of SAM (with ViT-B) and our deep instruction tuning methods on RefCOCO (Yu et al., 2016), i.e., the end-to-end and layer-wise ones, respectively [^2] Three methods are trained using the frozen ViT-B encoder.. In SAM, the text prompts only interact with the visual tokens in the shallow mask decoder, which is insufficient for cross-modal fusion and text-guided segmentation. Our simple deep tuning methods can improve the performance by a large margin without significantly changing the architecture of SAM.
With the proposed DIT methods, we apply SAM to the popular text-guided Referring Image Segmentation (RIS) (Liu et al., 2023c) task, and term the new models as DIT-SAMs. Extensive experiments are conducted on three highly competitive RIS benchmarks, namely RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016) and RefCOCOg (Mao et al., 2016; Nagaraja et al., 2016). The experimental results not only show the great advantages of DIT-SAMs over the default SAM, e.g., up to 11.6% improvement on RefCOCO testA, but also confirm the competitiveness of DIT-SAM as a stand alone RIS model. Overall, we provide a fast and feasible solution to extend SAM’s modal-modal ability.
@@
  • We propose two simple yet effective deep instruction tuning (DIT) methods for SAM, which are end-to-end and layer-wise, respectively. These DIT methods pave a quick and feasible way for the multi-modal extension of SAM.
@@
  • The proposed DIT-SAMs not only greatly outperforms the default SAM on RIS benchmarks, but also yield strong competitiveness against existing SOTA methods of RIS.
2. Related Work
2.1. Image Segmentation
@@
Image segmentation (Ronneberger et al., 2015; Long et al., 2015; He et al., 2017; Badrinarayanan et al., 2017) is a fundamental task in multimedia and computer vision, which requires pixel-level understanding of a given image. Recent years have witnessed its fast development supported by a bunch of segmentation-related tasks, such as semantic segmentation (Long et al., 2015; Badrinarayanan et al., 2017; Chen et al., 2017; He et al., 2017; Xie et al., 2021) which categorizes individual pixels into a predetermined set of classes, instance segmentation which primarily concerns the recognition and delineation of distinct object instances (He et al., 2017; Tian et al., 2020; Wang et al., 2020b), and panoptic segmentation which merges semantic and instance segmentation tasks by assigning class labels and instance identification (Kirillov et al., 2019; Li et al., 2019). In addition, a set of advanced segmentation networks (He et al., 2017; Tian et al., 2020; Wang et al., 2020b; Ma et al., 2021; Zu et al., 2023) are also proposed to promote the development of image segmentation. For instance, He et al. (He et al., 2017) propose Mask-RCNN for object instance segmentation. SOLO (Wang et al., 2020a) converts instance segmentation into a single-shot classification-solvable problem. IFR (Ma et al., 2021) is a implicit feature refinement module for high-quality instance segmentation, and TAR (Zu et al., 2023) presents the first attempt of weakly-supervised text instance segmentation by text recognition and segmentation. Recently, the Segment Anything project (Kirillov et al., 2023) proposes a powerful and general model for image segmentation, termed segment anything model (SAM). This project includes the creation of an unprecedentedly large segmentation dataset, so that SAM is trained with about 1 billion image-mask pairs. Via large-scale pre-training, SAM exhibits robust and versatile performance across various segmentation challenges which can also handle different types of prompts for conditional segmentation, such as point, box and text. More recently, SAM has been widely introduced to a set of segmentation-related tasks, such as low-level structural segmentation (Chen et al., 2023), surgical scene segmentation (Paranjape et al., 2023), and medical image segmentation (Guo et al., 2024; Xie et al., 2024; Quan et al., 2024). However, these applications mainly make use of the interactive capability of SAM on point or box prompts, while the text-instructed application is barely explored. Moreover, SAM also works much worse with texts than the other types of prompts due to the complexity of text information and the insufficient fusion of its default shallow decoder. In this paper, we propose deep instruction tuning for the multi-modal enhancement of SAM.
2.2. Referring Image Segmentation
@@
Referring image segmentation (RIS) (Yu et al., 2016) is one of the most popular text-conditioned segmentation task (Luo et al., 2020b, a; Jiao et al., 2021; Zhu et al., 2022; Shi et al., 2023; Liu et al., 2023f), which grounds the referent with binary masks according to the given natural language expression. Early research in RIS (Yu et al., 2016; Liu et al., 2019) typically adopts a two-stage pipeline and frames RIS as a region-text matching problem. One-stage RIS models (Liu et al., 2017; Li et al., 2018; Margffoy-Tuay et al., 2018) recently garner growing attention from both academia and industry. In these one-stage models, text features are often embedded into the vision network and directly output the mask of the referent. Inspired by the success of Transformer, various Transformer-based RIS models (Zhu et al., 2022; Hu et al., 2023; Liu et al., 2023d, a; Wu et al., 2023c; Xu et al., 2023b) are recently proposed to improve the multi-modal reasoning ability for RIS. VLT (Ding et al., 2022) and SeqTR (Zhu et al., 2022) use a encoder-decoder framework to fuse visual and language inputs into the decoder. ReSTR (Kim et al., 2022) employs two separate Transformer encoders to handle two different modalities. More recently, attempts have also been made to explore a wider range of text conditional segmentation by combining large language models (LLMs) and training with large-scale image-text pairs (Liu et al., 2023g; Yan et al., 2023; Lai et al., 2023; Zou et al., 2023). For instance, UniLSeg (Liu et al., 2023g), LISA (Lai et al., 2023) and SEEM (Zou et al., 2023) focus on text-instructed open-world segmentation, which typically require massive multi-modal training data and LLMs like LLaMA-7B (Touvron et al., 2023). Compared with these advancements, we mainly focus on addressing the text awareness of SAM and only conduct quick validations on RIS benchmarks. Among these progresses, the most relevant works to our DITs are LAVT (Yang et al., 2022) and ETRIS (Li and Sigal, 2021), which also only use two modality encoders for deep multi-modal fusion. However, our DITs greatly differ in both principle and designs. In particular, LAVT and ETRIS aim to achieve multi-modal interactions via cross-attention-based modules between two Transformers. In contrast, the principle of DITs is to project the text words onto the semantic space of SAM’s image encoder, thus achieving feature fusion and learning in one forward network. This property also makes DIT’s designs more intuitve and simpler, which also exhibits obvious merits in both training costs and performance.
@@
Figure 3. Overall pipeline of Deep Instruction Tuning (DIT). We propose (a) End-to-end DIT (E-DIT) that appends language instructions before visual features to achieve early-fusion for enhancing multi-modal reasoning capability; (b) Layer-wise DIT (L-DIT) that layer-wisely projects to text tokens interact with the existing sequences in the well learned semantic space.
2.3. Prompt Tuning
With the ever increasing parameter size of pre-trained models (Devlin et al., 2018; Dosovitskiy et al., 2020; Touvron et al., 2023), the direct fine-tune on downstream tasks becomes prohibitively expensive. In this case, prompt tuning (Gu et al., 2021; Liu et al., 2021a; Wei et al., 2021; Han et al., 2022) is proposed to use hand-craft or learnable prompt tokens to mitigate distribution shifts between pre-training and downstream tasks. The significant achievements of prompt tuning in natural language processing (NLP) have also inspired its application to computer vision and vision-language research (Zhou et al., 2022b, a; Wu et al., 2023b; Luo et al., 2023; Liu et al., 2023e). CoOp (Zhou et al., 2022b) remains CLIP parameters unchanged and uses learnable vectors as soft prompts for the input text, thereby maximizing CLIP’s zero-shot retrieval capabilities. To address the overfitting issue on base classes, CoCoOp (Zhou et al., 2022a) proposes instance adaptation by incorporating the visual features of each image into the parameterized prompts. RPO (Lee et al., 2023) enhances the robustness of VL models by masked attention to prevent the added learnable embeddings from altering the original internal representations of the model. More recently, Wu et al. (Wu et al., 2023b) propose a method of approximated prompt tuning to reduce the additional cost of the inserted prompt tokens. Regarding soft prompt tuning (Lester et al., 2021; Luo et al., 2023; Wu et al., 2023d), whether applied deeply or shallowly, it employs learnable tokens to assist the model in adapting to downstream data distributions without the necessity for complete tuning. Although the proposed DIT methods and prompt tuning are quite similar in process, but they still hold different principles and designs. To explain, prompt tuning often adopts learnable or hand-craft tokens to brdige the gap between pre-training and downstream tasks, which are often static to tasks (Jia et al., 2022). In contrast, DITs aim to project the text tokens onto the semantic space of SAM, which are dynamic for different examples. Conclusively, prompt tuning primarily facilitates rapid task adaptation, whereas our DITs offer a novel approach to VL reasoning for SAM.
3. Method
3.1. Problem Definition
@@
Referring image segmentation (RIS) (Yu et al., 2016) aims to segment the referred object in an image $I$ conditioned on the paired natural language expression $T$, its objective is formulated as
(1) $\arg\min_{\theta}\mathcal{L}(\mathcal{F}(I,T;\theta),M),$
@@
To solve this task, SAM (Kirillov et al., 2023) first extracts visual and text features through its vision and language encoders, based on which the light-weight decoder takes as input to predict the binary masks. Thus, the model $\mathcal{F}$ instantiated by SAM is
(2) $\mathcal{F}(I,T;\theta)=\mathcal{F}_{d}\big{(}\mathcal{F}_{v}(I),\mathcal{F}_{t}(T)\big{)},$
@@
where $\mathcal{F}_{v}$, $\mathcal{F}_{t}$ and $\mathcal{F}_{d}$ refer to visual backbone, language encoder and mask decoder, respectively.
@@
The mask decoder is very lightweight and only contains two cross-attention layers, which are in charge of both mask decoding and multi-modal interactions. In this case, SAM can be regarded as one of the late-fusion models which proved to be less effective than the early-fusion ones (Luo et al., 2020a, b; Zhu et al., 2022; Liu et al., 2023c) in addressing the complicated multi-modal reasoning tasks. Meanwhile, due to the lack of training on massive cross-modal data, text-guided SAM performs worse than the point- or box-guided SAM, leading to much inferior cross-modal abilities than existing SOTA methods in RIS (Zhu et al., 2022; Liu et al., 2023c).
@@
We introduce Deep Instruction Tuning (DIT) for SAM to better follow the intentions reflected in the natural language expression $T$. In particular, DIT keeps the whole SAM architecture unchanged and directly uses the image encoder as a strong multi-modal fusion network. The text instruction words can continually interact with the visual contents through the whole feature learning process, helping SAM progressively locate the referent. Specifically, SAM with DIT can be then formulated as
(3) $\mathcal{F}(T,I;\theta)=\mathcal{F}_{d}\big{(}\mathcal{F}_{v}(I,\phi(\mathcal{F}_{t}(T))),\mathcal{F}_{t}(T)\big{)}.$
@@
Here, $\phi(\cdot)$ is a mapping function to project the text features onto the default visual space of SAM.
3.2. End-to-end Deep Instruction Tuning